let F(u,v) where u=u(x,y,z) and v=v(x,y,z), further let z=z(x,y). show that on elimination of the arbitrary function F we obtain the first order partial differential equation;

Pp+Qq=R


1
Expert's answer
2021-10-25T14:55:26-0400

Clearly "F(u, v) = 0" is an implicit function.


"0=F_x=\dfrac{\partial F }{\partial u}\dfrac{\partial u}{\partial x}+\dfrac{\partial F }{\partial v}\dfrac{\partial v}{\partial x}"

"0=F_y=\dfrac{\partial F }{\partial u}\dfrac{\partial u}{\partial y}+\dfrac{\partial F }{\partial v}\dfrac{\partial v}{\partial y}"

"\dfrac{\partial F }{\partial v}=-\dfrac{\dfrac{\partial F }{\partial u}\dfrac{\partial u}{\partial y}}{\dfrac{\partial v}{\partial y}}"

Substitute


"\dfrac{\partial F }{\partial u}\dfrac{\partial u}{\partial x}-\dfrac{\dfrac{\partial F }{\partial u}\dfrac{\partial u}{\partial y}}{\dfrac{\partial v}{\partial y}}\dfrac{\partial v}{\partial x}=0"

"\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}=0"

"\dfrac{\partial u}{\partial x}=\dfrac{\partial u}{\partial x}+\dfrac{\partial u}{\partial z}\dfrac{\partial z}{\partial x}"

"\dfrac{\partial v}{\partial x}=\dfrac{\partial v}{\partial x}+\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial x}"

"\dfrac{\partial u}{\partial y}=\dfrac{\partial u}{\partial y}+\dfrac{\partial u}{\partial z}\dfrac{\partial z}{\partial y}"

"\dfrac{\partial v}{\partial y}=\dfrac{\partial v}{\partial y}+\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial y}"

Then


"(\dfrac{\partial u}{\partial x}+\dfrac{\partial u}{\partial z}\dfrac{\partial z}{\partial x})(\dfrac{\partial v}{\partial y}+\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial y})"

"-(\dfrac{\partial u}{\partial y}+\dfrac{\partial u}{\partial z}\dfrac{\partial z}{\partial y})(\dfrac{\partial v}{\partial x}+\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial x})=0"



"\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}+\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial y}\dfrac{\partial z}{\partial x}+\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial y}+\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial x}\dfrac{\partial z}{\partial y}"

"-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial x}\dfrac{\partial z}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial x}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial z}\dfrac{\partial z}{\partial x}\dfrac{\partial z}{\partial y}=0"


"(\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial z}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial y})\dfrac{\partial z}{\partial x}+(\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial x}-\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial z})\dfrac{\partial z}{\partial y}"

"=\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}"

"P=\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial z}-\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial y}"

"Q=\dfrac{\partial u}{\partial z}\dfrac{\partial v}{\partial x}-\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial z}"

"R=\dfrac{\partial u}{\partial x}\dfrac{\partial v}{\partial y}-\dfrac{\partial u}{\partial y}\dfrac{\partial v}{\partial x}"

"p=\dfrac{\partial z}{\partial x}, q=\dfrac{\partial z}{\partial y}"

"Pp+Qq=R"